OpenStudy (anonymous):

integral of square root of (1-cos x)dx

OpenStudy (anonymous):

\[\int\limits_{?}^{?} \sqrt{1-\cos x} dx\] is this the question?

OpenStudy (mr.math):

Use a substitution \(u=1-\cos{x}\). I think it will work.

OpenStudy (anonymous):

yes, and no a u sub won't work du = sinx dx, which we don't have

OpenStudy (anonymous):

yes to janie no to mr. math

OpenStudy (anonymous):

it is a crazy integral, and I know there is a way to solve it, I just don't know how to get at it.

OpenStudy (anonymous):

try wolframalfpha

OpenStudy (anonymous):

man, I just couldn't get an answer with the steps shown when I googled wolfram, thanks a bunch janie

OpenStudy (anonymous):

oops, nevermind. I forgot the sqrt. you should just ask your teacher - he/she would do a much better job of explaining this problem than me.

OpenStudy (anonymous):

How about a trig half angle substitution....if you need more hints letm me know

OpenStudy (anonymous):

this one is tricky - i just feel that u = 1-cosx should work - i'll try again

OpenStudy (anonymous):

use the substitution for sin(x/2)

OpenStudy (anonymous):

it will work out nice.

OpenStudy (mr.math):

I'm sure the substitution will work. Here is a hint: \[1-\cos{x}=u \implies \cos{x}=1-u \implies \sin{x}=\sqrt{2u-u^2}.\]

OpenStudy (mr.math):

You can draw a triangle to get that.

OpenStudy (mr.math):

So \[du=\sin{x}dx \implies dx=\frac{du}{\sqrt{2u-u^2}}.\] If we call our original integral I, then \[I=\int \frac{\sqrt{u}}{\sqrt{u}\sqrt{2-u}}du=\int \frac{du}{\sqrt{2-u}}.\]

OpenStudy (mr.math):

Which is a straight forward integral.

OpenStudy (anonymous):

right math - thats the way to go

OpenStudy (anonymous):

now use substitution t = 2-u

OpenStudy (anonymous):

integral - dt/sqrt t

OpenStudy (anonymous):

sorry thats dt/sqrt t integral of this is 2 sqrt t = 2 sqrt (2-u) = 2 sqrt (cosx + 1) + c

OpenStudy (nikvist):

\[\cos{x}=\cos^2{x/2}-\sin^2{x/2}=1-\sin^2{x/2}-\sin^2{x/2}=1-2\sin^2{x/2}\]\[\Rightarrow 1-\cos{x}=2\sin^2{x/2}\]

OpenStudy (anonymous):

 

OpenStudy (anonymous):

thanks a ton guys, I had to run out but Mr. Math was totally right on.

OpenStudy (anonymous):

pythagorean correlation I didn't take into account for the sin(x)dx, wonderful stuff.